# An Asymptotically Optimal Coordinate Descent Algorithm for Learning Bayesian Networks from Gaussian Models

### v1

@@
Abstract.
This paper studies the problem of learning Bayesian networks from continuous observational data, generated according to a linear Gaussian structural equation model. We consider an $\ell_{0}$-penalized maximum likelihood estimator for this problem which is known to have favorable statistical properties but is computationally challenging to solve, especially for medium-sized Bayesian networks. We propose a new coordinate descent algorithm to approximate this estimator and prove several remarkable properties of our procedure: the algorithm converges to a coordinate-wise minimum, and despite the non-convexity of the loss function, as the sample size tends to infinity, the objective value of the coordinate descent solution converges to the optimal objective value of the $\ell_{0}$-penalized maximum likelihood estimator. Finite-sample optimality and statistical consistency guarantees are also established. To the best of our knowledge, our proposal is the first coordinate descent procedure endowed with optimality and statistical guarantees in the context of learning Bayesian networks. Numerical experiments on synthetic and real data demonstrate that our coordinate descent method can obtain near-optimal solutions while being scalable.
1. Introduction
1.1. Background and Related Work
@@
Three broad classes of methods for learning DAGs from data are constraint-based, score-based, and hybrid. Constraint-based methods use repeated conditional independence tests to determine the presence of edges in a DAG. A prominent example is the PC algorithm and its extensions [20, 21]. While the PC algorithm can be applied in non-parametric settings, testing for conditional independencies is generally hard [17]. Furthermore, even in the Gaussian setting, statistical consistency guarantees for the PC algorithm are shown under the strong faithfulness condition [12], which is known to be restrictive in high-dimensional settings [22]. Score-based methods often deploy a penalized log-likelihood as a score function and search over the space of DAGs to identify a DAG with an optimal score. These approaches do not require the strong faithfulness assumption. However, statistical guarantees are not provided for many score-based approaches and solving them exactly suffers from high computational complexity. For example, learning an optimal graph using dynamic programming takes about 10 hours for a medium-size problem with 29 nodes [18]. Several papers [13, 25] offer speedup by casting the problem as a convex mixed-integer program, but finding an optimal solution with these approaches can still take an hour for a medium-sized problem. Finally, hybrid approaches combine constraint-based and score-based methods by using background knowledge or conditional independence tests to restrict the DAG search space [21, 16].
@@
Several strategies have been developed to make score-based methods more scalable by finding approximate solutions instead of finding optimally scoring DAGs. One direction to find good approximate solutions is to resort to greedy-based methods, with a prominent example being the Greedy Equivalence Search (GES) algorithm [5]. GES performs a greedy search on the space of completed partially directed acyclic graphs (an equivalence class of DAGs) and is known to produce asymptotically consistent solutions [5]. Despite its favorable properties, GES does not provide optimality or consistency guarantees for any finite sample size. Further, the guarantees of GES assume a fixed number of nodes with sample size going to infinity and do not allow for a growing number of nodes. Another direction is gradient-based approaches [27, 28], which relax the discrete search space over DAGs to a continuous search space, allowing gradient descent and other techniques from continuous optimization to be applied. However, the search space for these problems is highly non-convex, resulting in limited guarantees for convergence, even to a local minimum. Finally, another notable direction is based on coordinate descent; that is iteratively maximizing the given score function over a single parameter, while keeping the remaining parameters fixed and checking that the resulting model is a DAG at each update [1, 2, 9, 26]. While coordinate descent algorithms have shown significant promise in learning large-scale Bayesian networks, to the best of our knowledge, they do not come with convergence, optimality, and statistical guarantees.
1.2. Our Contributions
@@
We propose a new score-based coordinate descent algorithm for learning Bayesian networks from Gaussian linear structural equation models. Remarkably, unlike prior coordinate descent algorithms for learning Bayesian networks, our procedure provably i) converges to a coordinate-wise minimum, ii) produces optimally scoring DAGs as the sample size tends to infinity despite the non-convex nature of the problem, and iii) yields asymptotically consistent estimates that also provide finite-sample guarantees that allow for a growing number of nodes. As a scoring function for this approach, we deploy an $\ell_{0}$-penalized Gaussian log-likelihood, which implies that optimally-scoring DAGs are solutions to a highly non-convex $\ell_{0}$-penalized maximum likelihood estimator. This estimator is known to have strong statistical consistency guarantees [23], but solving it is, in general, intractable. Thus, our coordinate descent algorithm can be viewed as a scalable and efficient approach to finding approximate solutions to this estimator that are asymptotically optimal (i.e., match the optimal objective value of the $\ell_{0}$ penalized maximum-likelihood estimator as the sample size tends to infinity) and have finite-sample statistical consistency guarantees.
@@
2. Problem Setup
2.1. Preliminaries and Definitions
Consider an unknown DAG whose $m$ nodes correspond to observed random variables $X\in\mathbb{R}^{m}$. We denote the DAG by $\mathcal{G}^{\star}=(V,E^{\star})$ where $V=\left\{1,\ldots,m\right\}$ is the vertex set and $E^{\star}\subseteq V\times V$ is the directed edge set. We assume that the random variables $X$ satisfy the linear structural equation model (SEM):
(1) $X={B^{\star}}^{\mathrm{\scriptscriptstyle T}}X+\epsilon,$
where $B^{\star}\in\mathbb{R}^{m\times m}$ is the connectivity matrix with zeros on the diagonal and $B^{\star}_{jk}\neq 0$ if $(j,k)\in E^{\star}$. In other words, the sparsity pattern of $B^{\star}$ encodes the true DAG structure. Further, $\epsilon\sim\mathcal{N}(0,\Omega^{\star})$ is a random Gaussian noise vector with zero mean and independent coordinates so that $\Omega^{\star}$ is a diagonal matrix. Assuming, without loss of generality, that all random variables are centered, each variable $X_{j}$ in this model can be expressed as the linear combination of its parents—the set of nodes with directed edges pointing to $j$—plus independent Gaussian noise. By the SEM (1) and the Gaussianity of $\epsilon$, the random vector $X$ follows the Gaussian distribution $\mathcal{P}^{\star}=\mathcal{N}(0,\Sigma^{\star})$, with $\Sigma^{\star}=(I-B^{\star})^{\mathrm{\scriptscriptstyle-T}}\Omega^{\star}(I-B^{\star})^{\mathrm{\scriptscriptstyle-1}}$. Throughout, we assume that the distribution $\mathcal{P}^{\star}$ is non-degenerate, or equivalently, $\Sigma^{\star}$ is positive definite. Our objective is to estimate the matrix $B^{\star}$, or as we describe next, an equivalence class when the underlying model is not identifiable.
@@ 2.1. Preliminaries and Definitions
To see why the model (1) is generally not identifiable, note that there are multiple tuples $(B,\Omega)$ where $\mathcal{G}(B)$ is DAG and $\Omega$ is a positive definite diagonal matrix with $\Sigma^{\star}=(I-B)^{\mathrm{\scriptscriptstyle-T}}\Omega(I-B)^{\mathrm{\scriptscriptstyle-1}}$ [23]. As a result, the SEM given by $(B,\Omega)$ yields an equally representative model as the one given by the population parameters $(B^{\star},\Omega^{\star})$. When $\mathcal{G}^{\star}$ is faithful with respect to the graph $\mathcal{G}^{\star}$ (see Assumption 7 in Section 4 for a formal definition), the sparsest DAGs that are compatible with $\mathcal{P}^{\star}$ are precisely $\mathrm{MEC}(\mathcal{G}^{\star})$, the Markov equivalence class of $\mathcal{G}^{\star}$ [23]. Next, we formally define the Markov equivalence class.
Definition 2.
@@
(Markov equivalence class $\mathrm{MEC}(\mathcal{G})$[24]) Let $\mathcal{G}=(V,E)$ be a DAG. Then, $\mathrm{MEC}(\mathcal{G})$ consists of DAGs that have the same skeleton and same v-structures as $\mathcal{G}$. The skeleton of $\mathcal{G}$ is the undirected graph obtained from $\mathcal{G}$ by substituting directed edges with undirected ones. Furthermore, nodes $i,j$, and $k$ form a v-structure if $(i,k)\in E$ and $(j,k)\in E$, and there is no edge between $i$ and $j$.
2.2. $\ell_{0}$-Penalized Maximum Likelihood Estimator
@@
Consider $n$ independent and identically distributed observations of the random vector $X$ generated according to (1). Let $\hat{\Sigma}$ be the sample covariance matrix obtained from these observations. Further, consider a Gaussian SEM parameterized by connectivity matrix $B$ and noise variance $\Omega$ with $D=\Omega^{-1}$. The parameters $(B,D)$ specify the following precision, or inverse covariance, matrix $\Theta:=\Theta(B,D):=(I-B){D}(I-B)^{\mathrm{\scriptscriptstyle T}}$. The negative log-likelihood of this SEM is proportional to $\ell_{n}(\Theta)=\mathrm{trace}(\Theta\hat{\Sigma})-\log\det(\Theta)$. Naturally, we seek a model that not only has a small negative log-likelihood but is also specified by a sparse connectivity matrix containing few nonzero elements. Thus, we deploy the following $\ell_{0}$-penalized maximum likelihood estimator with a regularization parameter $\lambda\geq 0$:
@@
(2) $\min_{B\in\mathbb{R}^{m\times m},D\in\mathbb{D}^{m}_{++}}\ell_{n}\left(\left(I-B\right)D\left(I-B\right)^{\mathrm{\scriptscriptstyle T}}\right)+\lambda^{2}\|B\|_{\ell_{0}}\quad\text{s.t.}\quad\mathcal{G}(B)\text{ is a DAG}.$
@@
Here, $\mathbb{D}^{m}_{++}$ denotes the collection of positive definite $m\times m$ diagonal matrices and $\|B\|_{\ell_{0}}$ denotes the number of non-zeros in $B$. Note that the $\ell_{0}$ penalty is generally preferred over the $\ell_{1}$ penalty or minimax concave penalty (MCP) for penalizing the complexity of the model. In particular, $\ell_{0}$ regularization exhibits the important property that equivalent DAGs—those in the same Markov equivalence class—have the same penalized likelihood score, while this is not the case for $\ell_{1}$ or MCP regularization [23]. Indeed, this lack of score invariance with $\ell_{1}$ regularization partially explains the unfavorable properties of some existing methods (see Section 5).
@@
The Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{B}^{\mathrm{opt}}))$ of the connectivity matrix $\hat{B}^{\mathrm{opt}}$ obtained from solving (2) provides an estimate of $\mathrm{MEC}(\mathcal{G}^{\star})$. van de Geer and Bühlmann [23] prove that this estimate has desirable statistical properties; however, solving it is, in general, intractable. As stated, the objective function $\ell_{n}((I-B)D(I-B)^{\mathrm{\scriptscriptstyle T}})$ is non-convex and non-linear function of $(B,D)$. Furthermore, the $\log\det$ function in the likelihood $\ell_{n}$ is not amenable to standard mixed-integer programming optimization techniques. To circumvent the aforementioned challenges, Xu et al. [25] derive the following equivalent optimization model via the change of variables $\Gamma\leftarrow(I-B)D^{1/2}$:
@@
(3) $\min\limits_{\Gamma\in\mathbb{R}^{m\times m}}f(\Gamma)\quad\text{s.t.}\quad\mathcal{G}\left(\Gamma-\mathrm{diag}\left(\Gamma\right)\right)\text{ is a DAG}.$
@@
Here $f(\Gamma):=\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n})+\lambda^{2}\|\Gamma-\mathrm{diag}(\Gamma)\|_{\ell_{0}},$ and $\mathrm{diag}(\Gamma)$ is the diagonal matrix formed by taking the diagonal entries of $\Gamma$. The optimal solutions of (2) and (3) are directly connected: Letting $(\hat{B}^{\mathrm{opt}},\hat{D}^{\mathrm{opt}})$ be an optimal solution of (2), then $\hat{\Gamma}^{\mathrm{opt}}=(I-\hat{B}^{\mathrm{opt}})(\hat{D}^{\mathrm{opt}})^{1/2}$ is an optimal solution of (3). Furthermore, the sparsity pattern of $\hat{\Gamma}^{\mathrm{opt}}-\mathrm{diag}(\hat{\Gamma}^{\mathrm{opt}})$ is the same as that of $\hat{B}^{\mathrm{opt}}$; in other words, the Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{B}^{\mathrm{opt}}))$ is the same as the Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{\Gamma}^{\mathrm{opt}}-\mathrm{diag}(\hat{\Gamma}^{\mathrm{opt}})))$.
@@
Xu et al. [25] recast the optimization problem (3) as a convex mixed-integer program and provide algorithms to solve (3) to optimality. However, solving (3) is, in general, NP-hard, and obtaining optimality certificates may take an hour for a problem with 20 nodes [25].
3. A Coordinate Descent Algorithm for DAG Learning

### v3

@@
Abstract
This paper studies the problem of learning Bayesian networks from continuous observational data, generated according to a linear Gaussian structural equation model. We consider an $\ell_{0}$-penalized maximum likelihood estimator for this problem, which is known to have favorable statistical properties but is computationally challenging to solve, especially for medium-sized Bayesian networks. We propose a new coordinate descent algorithm to approximate this estimator and prove several remarkable properties of our procedure: The algorithm converges to a coordinate-wise minimum, and despite the non-convexity of the loss function, as the sample size tends to infinity, the objective value of the coordinate descent solution converges to the optimal objective value of the $\ell_{0}$-penalized maximum likelihood estimator. To the best of our knowledge, our proposal is the first coordinate descent procedure endowed with optimality guarantees in the context of learning Bayesian networks. Numerical experiments on synthetic and real data demonstrate that our coordinate descent method can obtain near-optimal solutions while being scalable.
Keywords: Directed acyclic graphs, $\ell_{0}$-penalization, Non-convex optimization, Structural equation models
1 Introduction
@@
Three broad classes of methods for learning DAGs from data are constraint-based, score-based, and hybrid. Constraint-based methods use repeated conditional independence tests to determine the presence of edges in a DAG. A prominent example is the PC algorithm and its extensions (Spirtes et al., 1993; Tsamardinos et al., 2006). While the PC algorithm can be applied in non-parametric settings, testing for conditional independencies is generally hard (Shah and Peters, 2018). Furthermore, even in the Gaussian setting, statistical consistency guarantees for the PC algorithm are shown under the strong faithfulness condition (Kalisch and Bühlmann, 2007), which is known to be restrictive in high-dimensional settings (Uhler et al., 2012). Score-based methods often deploy a penalized log-likelihood as a score function and search over the space of DAGs to identify a DAG with an optimal score. These approaches do not require the strong faithfulness assumption. However, statistical guarantees are not provided for many score-based approaches and solving them exactly suffers from high computational complexity. For example, learning an optimal graph using dynamic programming takes about 10 hours for a medium-size problem with 29 nodes (Silander and Myllymäki, 2006). Several papers (Küçükyavuz et al., 2023; Xu et al., 2024) offer speedup by casting the problem as a convex mixed-integer program, but finding an optimal solution with these approaches can still take an hour for a medium-sized problem. Finally, hybrid approaches combine constraint-based and score-based methods by using background knowledge or conditional independence tests to restrict the DAG search space (Tsamardinos et al., 2006; Nandy et al., 2018).
@@
Several strategies have been developed to make score-based methods more scalable by finding approximate solutions instead of finding optimally scoring DAGs. One direction to find good approximate solutions is to resort to greedy-based methods, with a prominent example being the Greedy Equivalence Search (GES) algorithm (Chickering, 2002). GES performs a greedy search on the space of completed partially directed acyclic graphs (an equivalence class of DAGs) and is known to produce asymptotically consistent solutions (Chickering, 2002). Despite its favorable properties, GES does not provide optimality or consistency guarantees for any finite sample size. Further, the guarantees of GES assume a fixed number of nodes with sample size going to infinity and do not allow for a growing number of nodes. Another direction is gradient-based approaches (Yu et al., 2019; Zheng et al., 2018), which relax the discrete search space over DAGs to a continuous search space, allowing gradient descent and other techniques from continuous optimization to be applied. However, the search space for these problems is highly non-convex, resulting in limited guarantees for convergence, even to a local minimum. Finally, another notable direction is based on coordinate descent; that is iteratively maximizing the given score function over a single parameter, while keeping the remaining parameters fixed and checking that the resulting model is a DAG at each update (Aragam and Zhou, 2015; Aragam et al., 2019; Fu and Zhou, 2013; Ye et al., 2020). While coordinate descent algorithms have shown significant promise in learning large-scale Bayesian networks, to the best of our knowledge, they do not come with convergence and optimality guarantees.
Contributions: We propose a new score-based coordinate descent algorithm for learning Bayesian networks from Gaussian linear structural equation models. Remarkably, unlike prior coordinate descent algorithms for learning Bayesian networks, our procedure provably i) converges to a coordinate-wise minimum, and ii) yields optimally scoring DAGs in the large sample limit despite the non-convex nature of the problem. Moreover, we characterize the convergence rate as a function of both the sample size and the number of nodes. A byproduct of our analysis is a characterization of the optimization landscape: in the large-sample regime, all local minima achieve objective values close a global minimum, and exactly match it in the limit—a result that may be of independent interest.
@@
As a scoring function for this approach, we deploy an $\ell_{0}$-penalized Gaussian log-likelihood, which implies that optimally-scoring DAGs are solutions to a highly non-convex $\ell_{0}$-penalized maximum likelihood estimator. This estimator is known to have strong statistical consistency guarantees (van de Geer and Bühlmann, 2013), but solving it is, in general, intractable. Thus, our coordinate descent algorithm can be viewed as a scalable and efficient approach to finding approximate solutions to this estimator that are asymptotically optimal (that is, match the optimal objective value of the $\ell_{0}$ penalized maximum-likelihood estimator as the sample size tends to infinity).
@@
2 Problem Setup
Consider an unknown DAG whose $m$ nodes correspond to observed random variables $X\in\mathbb{R}^{m}$. We denote the DAG by $\mathcal{G}^{\star}=(V,E^{\star})$ where $V=\left\{1,\ldots,m\right\}$ is the vertex set and $E^{\star}\subseteq V\times V$ is the directed edge set, where $(i,j)\in E^{\star}$ indicates $i\to j$ in graph $\mathcal{G}^{\star}$. We assume that the random variables $X$ satisfy the linear structural equation model (SEM):
$X={B^{\star}}^{\mathrm{\scriptscriptstyle T}}X+\epsilon,$ (1)
where $B^{\star}\in\mathbb{R}^{m\times m}$ is the connectivity matrix with zeros on the diagonal and $B^{\star}_{jk}\neq 0$ if $(j,k)\in E^{\star}$. In other words, the sparsity pattern of $B^{\star}$ encodes the true DAG structure. Further, $\epsilon\sim\mathcal{N}(0,\Omega^{\star})$ is a random Gaussian noise vector with zero mean and independent coordinates so that $\Omega^{\star}$ is a diagonal matrix. We use the Bachman-Landau symbols $\mathcal{O}$ to describe the limiting behavior of a function. Furthermore, we denote $z\asymp 1$ to express $z=\mathcal{O}(1)$ and $1/z=\mathcal{O}(1)$. Assuming, without loss of generality, that all random variables are centered, each variable $X_{j}$ in this model can be expressed as the linear combination of its parents—the set of nodes with directed edges pointing to jj—plus independent Gaussian noise. By the SEM (1) and the Gaussianity of $\epsilon$, the random vector $X$ follows the Gaussian distribution $\mathcal{P}^{\star}=\mathcal{N}(0,\Sigma^{\star})$, with $\Sigma^{\star}=(I-B^{\star})^{\mathrm{\scriptscriptstyle-T}}\Omega^{\star}(I-B^{\star})^{\mathrm{\scriptscriptstyle-1}}$. Throughout, we assume that the distribution $\mathcal{P}^{\star}$ is non-degenerate, or equivalently, $\Sigma^{\star}$ is positive definite. Our objective is to estimate the matrix $B^{\star}$, or as we describe next, an equivalence class when the underlying model is not identifiable.
@@ 2 Problem Setup
To see why the model (1) is generally not identifiable, note that there are multiple tuples $(B,\Omega)$ where $\mathcal{G}(B)$ is DAG and $\Omega$ is a positive definite diagonal matrix with $\Sigma^{\star}=(I-B)^{\mathrm{\scriptscriptstyle-T}}\Omega(I-B)^{\mathrm{\scriptscriptstyle-1}}$ (van de Geer and Bühlmann, 2013). As a result, the SEM given by $(B,\Omega)$ yields an equally representative model as the one given by the population parameters $(B^{\star},\Omega^{\star})$. When $\mathcal{G}^{\star}$ is faithful with respect to the graph $\mathcal{G}^{\star}$, the sparsest DAGs that are compatible with $\mathcal{P}^{\star}$ are precisely $\mathrm{MEC}(\mathcal{G}^{\star})$, the Markov equivalence class of $\mathcal{G}^{\star}$ (van de Geer and Bühlmann, 2013). Next, we formally define the Markov equivalence class.
Definition 2
@@
(Markov equivalence class $\mathrm{MEC}(\mathcal{G})$(Verma and Pearl, 1990)) Let $\mathcal{G}=(V,E)$ be a DAG. Then, $\mathrm{MEC}(\mathcal{G})$ consists of DAGs that have the same skeleton and same v-structures as $\mathcal{G}$. The skeleton of $\mathcal{G}$ is the undirected graph obtained from $\mathcal{G}$ by substituting directed edges with undirected ones. Furthermore, nodes $i,j$, and $k$ form a v-structure if $(i,k)\in E$ and $(j,k)\in E$, and there is no edge between $i$ and $j$.
@@
With these preliminaries and definitions in place, we next describe the proposed $\ell_{0}$-penalized maximum likelihood estimator. Consider $n$ independent and identically distributed observations of the random vector $X$ generated according to (1). Let $\hat{\Sigma}$ be the sample covariance matrix obtained from these observations. Further, consider a Gaussian SEM parameterized by connectivity matrix $B$ and noise variance $\Omega$ with $D=\Omega^{-1}$. The parameters $(B,D)$ specify the following precision, or inverse covariance, matrix $\Theta:=\Theta(B,D):=(I-B){D}(I-B)^{\mathrm{\scriptscriptstyle T}}$. The negative log-likelihood of this SEM is proportional to $\ell_{n}(\Theta)=\mathrm{trace}(\Theta\hat{\Sigma})-\log\det(\Theta)$. Naturally, we seek a model that not only has a small negative log-likelihood but is also specified by a sparse connectivity matrix containing few nonzero elements. Thus, we deploy the following $\ell_{0}$-penalized maximum likelihood estimator with a regularization parameter $\lambda\geq 0$:
@@
$\min_{B\in\mathbb{R}^{m\times m},D\in\mathbb{D}^{m}_{++}}\ell_{n}\left(\left(I-B\right)D\left(I-B\right)^{\mathrm{\scriptscriptstyle T}}\right)+\lambda^{2}\|B\|_{\ell_{0}}\quad\text{s.t.}\quad\mathcal{G}(B)\text{ is a DAG}.$ (2)
@@
Here, $\mathbb{D}^{m}_{++}$ denotes the collection of positive definite $m\times m$ diagonal matrices and $\|B\|_{\ell_{0}}$ denotes the number of non-zeros in $B$. Note that the $\ell_{0}$ penalty is generally preferred over the $\ell_{1}$ penalty or minimax concave penalty (MCP) for penalizing the complexity of the model. In particular, $\ell_{0}$ regularization exhibits the important property that equivalent DAGs—those in the same Markov equivalence class—have the same penalized likelihood score, while this is not the case for $\ell_{1}$ or MCP regularization (van de Geer and Bühlmann, 2013). Indeed, this lack of score invariance with $\ell_{1}$ regularization partially explains the unfavorable properties of some existing methods (see Section 6).
@@
The Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{B}^{\mathrm{opt}}))$ of the connectivity matrix $\hat{B}^{\mathrm{opt}}$ obtained from solving (2) provides an estimate of $\mathrm{MEC}(\mathcal{G}^{\star})$. van de Geer and Bühlmann (2013) prove that this estimate has desirable statistical properties; however, solving it is, in general, intractable. As stated, the objective function $\ell_{n}((I-B)D(I-B)^{\mathrm{\scriptscriptstyle T}})$ is non-convex and non-linear function of $(B,D)$. Furthermore, the $\log\det$ function in the likelihood $\ell_{n}$ is not amenable to standard mixed-integer programming optimization techniques. To circumvent the aforementioned challenges, Xu et al. (2024) derive the following equivalent optimization model via the change of variables $\Gamma\leftarrow(I-B)D^{1/2}$:
@@
$\min\limits_{\Gamma\in\mathbb{R}^{m\times m}}f(\Gamma)\quad\text{s.t.}\quad\mathcal{G}\left(\Gamma-\mathrm{diag}\left(\Gamma\right)\right)\text{ is a DAG}.$ (3)
@@
Here $f(\Gamma):=\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})+\lambda^{2}\|\Gamma-\mathrm{diag}(\Gamma)\|_{\ell_{0}},$ and $\mathrm{diag}(\Gamma)$ is the diagonal matrix formed by taking the diagonal entries of $\Gamma$. The optimal solutions of (2) and (3) are directly connected: Letting $(\hat{B}^{\mathrm{opt}},\hat{D}^{\mathrm{opt}})$ be an optimal solution of (2), then $\hat{\Gamma}^{\mathrm{opt}}=(I-\hat{B}^{\mathrm{opt}})(\hat{D}^{\mathrm{opt}})^{1/2}$ is an optimal solution of (3). Furthermore, the sparsity pattern of $\hat{\Gamma}^{\mathrm{opt}}-\mathrm{diag}(\hat{\Gamma}^{\mathrm{opt}})$ is the same as that of $\hat{B}^{\mathrm{opt}}$; in other words, the Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{B}^{\mathrm{opt}}))$ is the same as the Markov equivalence class $\mathrm{MEC}(\mathcal{G}(\hat{\Gamma}^{\mathrm{opt}}-\mathrm{diag}(\hat{\Gamma}^{\mathrm{opt}})))$.
@@
Xu et al. (2024) recast the optimization problem (3) as a convex mixed-integer program and provide algorithms to solve (3) to optimality. However, solving (3) is, in general, NP-hard, and obtaining optimality certificates may take an hour for a problem with 20 nodes (Xu et al., 2024).
3 A Coordinate Descent Algorithm for DAG Learning